• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R Roche

    Log entry time 00:00:19 on August 27,1999

    Entry number 26263

    keyword=Shift summary

    Shift Summary: 16:00-midnight

    16:00 Moller measurement in progress.
    17:45 Moller finished. Ready to take hydrogen data.
    17:50 Problems with beam due to storm.
    18:03 Beam back. Beginning run #1467.
    18:55 Begin run #1468.
    19:03 Stevens tells us that the temperature in the 9th dipole room is too low (74 vas 90); does not matter as long as we do not make an arc measurement. CFP
    19:57 Begin run #1469. Beam current dropped from 54.5 to 50 uA during run.
    20:57 Begin run #1470. Beam current back up to 55.4 uA.
    22:00 start run 1471
    22:03 reset HV FPP 8.1.18. May have been off for last 1.5 hr. Did not want to reset while taking data.
    22:35 reset HV FPP 8.1.18. Was off only about 10 min, data taking not interrupted.
    23:00 Begin run #1472.
    23:26 End run #1472. Junk. High voltage trip. Begin run #1473.
    23:54 Finished 9.3 M hydrogen events. Switching to carbon target.